OpenStudy (anonymous):

How would I solve this? Create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number.

OpenStudy (math&ing001):

Create a random function like h(x) = x^2 and f(x) = x+1 When you plug in a number you'll get different results for f(h(x)) and h(f(x)).

OpenStudy (anonymous):

Thanks :)

OpenStudy (math&ing001):

Welcome !